OpenStudy (sarahezzmck):

Can someone help me with this please? Solve for x: 5 over quantity x squared minus 4 plus 2 over x equals 2 over quantity x minus 2

OpenStudy (kewlgeek555):

\[\frac{ 5 }{ x ^{2} }-4+\frac{ 2 }{ x }-2\]I hate it when they put equations as words! But I believe this is the correct equation.

OpenStudy (sarahezzmck):

I'm sorry, :( I forgot about putting them into the actual equation, but I actually just got the answer to that one! Could you possibly answer another question of mine please?

OpenStudy (kewlgeek555):

I could try. ;}

OpenStudy (sarahezzmck):

Thank you! And btw, go to the same school, OCVS. lol. So the question is Solve for X: \[\frac{ 7 }{ 8 }-\frac{ 1 }{ x}\]

OpenStudy (sarahezzmck):

\[=\frac{ 3 }{ 4 }\]

OpenStudy (kewlgeek555):

You go to OCVS? What grade are you in? Anyways, x = 8.

OpenStudy (sarahezzmck):

10th. Lol, but i'm terrible at algebra, I'm good at everything else, like English, History and Biology but Algebra....total FAIL. I have Hogan for English, Fields for History and Besuden for Biology and then Schonborn for Algebra

OpenStudy (sarahezzmck):

what grade are you in? and also I have one more question :D

OpenStudy (kewlgeek555):

10th? I'm in ninth grade. Hm, did you go to the Courthouse field trip - if you did, then you could've seen me!

OpenStudy (sarahezzmck):

no, there were no more spaces and I wasn't allowed. I'm not allowed to go to any of the workdays anymore either lol. We have a joke, in my "group" at the workdays, we say, "Workdays! Where you go to get NO work done" because one theres too many kids, and two everyones talking lol.

OpenStudy (kewlgeek555):

You're not allowed to go to the workdays? I was looking forward to meeting you!

OpenStudy (sarahezzmck):

So the second question I have for you is, \[\frac{ 2 }{ x+2 }+\frac{ 1 }{ 5}=\frac{ 6 }{ x+5 } solve for x. \]

OpenStudy (sarahezzmck):

haha, after yesterday, noooo. I got alll my workdone but i failed most of it. :(

OpenStudy (sarahezzmck):

that was my "last chance"

OpenStudy (kewlgeek555):

I got 0, 13. Anyways, so that's why? Hm, they must be getting strict.

OpenStudy (sarahezzmck):

No, my family is. But do you know michael Rodgers? he's in 9th. And thank you btw :) i appreciate it

OpenStudy (sarahezzmck):

OMG i love you I just got 10 out of 10!!! lol

OpenStudy (kewlgeek555):

Well, I guess I don't know because I am rather shy and to tell you the truth, I don't do work on Wordays, unless the teacher is coming up to me. But I have good grades. I do not know Michael Rogers. Should I meet him next time?

OpenStudy (sarahezzmck):

lol, yeah. He's a good buddy of mine and really friendly. He was one of the first people I met there. So could you help me with this last quiz I have to take? Please, that is. :)

OpenStudy (sarahezzmck):

You're like my lifesaver right now bro! :)

OpenStudy (kewlgeek555):

Yes, I can try to help you with the last quiz.

OpenStudy (sarahezzmck):

Thank you :) what is the discontinuity of the function f(x)=\[x^{2}-3x-28 \over x+4\]

OpenStudy (kewlgeek555):

o.o I know that solving it would be x-7, but since I am not in Algebra II, yet, I do not know what is a "discontinuity".

OpenStudy (sarahezzmck):

Idk either. I hate that they do stuff like that and then I go back into the lessona dnit says NOTHING about that!. Lol, thanks anyways :)

OpenStudy (kewlgeek555):

Oh well. I think the new Common Core Standards ruined everything.

OpenStudy (sarahezzmck):

But they are like poitn on a graph though, btw for your future :)

OpenStudy (kewlgeek555):

Oh, okay.

OpenStudy (sarahezzmck):

lol whys that?

OpenStudy (kewlgeek555):

thanks.

OpenStudy (kewlgeek555):

It just made everything harder.

OpenStudy (kewlgeek555):

They don't know how to explain, and even the teachers are complaining! Even for little kids - they teach in younger grades with "counters" instead of numbers.

OpenStudy (sarahezzmck):

Shurwelcome :) oh yeah thats true. so another question, if you know i, yay, if not, its alll good. :) What are the zero(s) of the function, f(x)= \[4x ^{2}-36x \over x-9\]

OpenStudy (sarahezzmck):

I just hate tests in general, including the EOCs and FCAT. Kids get so stressed out about, including myself, and bc there so stressed they fail it. I do terribly on tests because I stress out and then freak out, forget everything I thought I learned and fail...miserably.

OpenStudy (kewlgeek555):

You know what if funny? I always get a B in Math and an A+ in English and I LOVE writing and last year I got a failing grade in writing, I got an way above average score in math and I got an average score in English.

OpenStudy (sarahezzmck):

lol, nice. So I'm gonna have to close this question bc I need help with a couple more questions but I think we can still talk on this one...im not sure though tbh.

OpenStudy (kewlgeek555):

We can.

OpenStudy (kewlgeek555):

Anyways, I got 0...

OpenStudy (sarahezzmck):

Thank you :) so how long have you been doing virtual school?

OpenStudy (kewlgeek555):

Since sixth grade.

OpenStudy (sarahezzmck):

oooh, this is my first year.. I went to Oak Ridge last year.

OpenStudy (sarahezzmck):

where the ghetto kids go, lmao except im not ghetto....im a white country girl who dreams of moving to Australia, thats why my pic on here is of Blue Lake in Australia

OpenStudy (kewlgeek555):

Oak Ridge High School? That's close to where I used to work. Anyways, since I am a coward, I say that if I go to Australia, it is for a one week vacation MAX and it will only be on a 5-star building on one of the top floors because I dont want any spiders or snakes tryng to kill me. I also wont go in the water because of octopus, stonefish, snakes, jellyfish, alligators, and sharks.